# Local-global compatibility and the exceptional zero conjecture for $\operatorname{GL}(3)$

### v1

@@ 9.6.2. Some important lemmas
where $0\leqslant y<z\leqslant n+1$ are integers such that j=(y+1)(n+1)(y+1)y/2zj=(y+1)(n+1)-(y+1)y/2-z (if $y=0$, then ${\mathcal{B}}_{j}=\{z\}$, and if $z=n+1$, then ${\mathcal{B}}_{j}=\{1,\dots,y\}$). We set ${\mathcal{B}}_{0}=\varnothing$. Let $\lambda=(\lambda_{1},\dots,\lambda_{n})$ be a $\operatorname{GL}_{n}$-dominant character, and let $a\in\mathbb{Z}$ be an integer. Set $x_{j}=w_{{\mathcal{B}}_{j}}$ and $w_{j}=w_{{\mathcal{B}}_{j}^{c}}$.
@@
works (i.e., $\lambda=(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{i}=(2^{n}-2^{i})M$).
@@ 9.6.3. Mixed characteristic local-global compatibility
$\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})\cdot m\cdot\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})]\mapsto\overline{\psi}(\operatorname{det}m)\cdot\Theta([\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})\cdot m\cdot\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})]),\quad\quad m\in\operatorname{GL}(\mathbb{A}_{f}^{S}),$
@@
We first note that a Galois representation $\varrho_{j}$ satisfying (1) already exists from the work of Scholze [Sch15]. We therefore want to verify (2) and (3) for this representation. Since this only concerns the behaviour at $p$, we are free to enlarge $S$, increase $\mathcal{O}$ and deepen the level $K$ at primes in $S-\{p\}$ as we wish. Here we are using the fact that Galois representations over the Henselian semi-local $\mathcal{O}$-algebra $\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c,\lambda;j)$, reducing to $\overline{\rho}_{\mathfrak{m}_{0}}$ modulo each of its maximal ideals, are determined (up to isomorphism) by their associated determinant (see [Che14, Theorem 2.22]).
@@
But $\operatorname{H}^{0}(X_{\operatorname{GL}_{1},C},\mathcal{O})$ a finite free $\mathcal{O}$-module with rank equal to the number of connected components of $X_{\operatorname{GL}_{1},C}$, and contains the trivial representation $\mathcal{O}$ of the Hecke algebra for $\operatorname{GL}_{1}$. This implies that we have a natural map
@@
induced from mapping a Hecke operator for $M$ to one for $\operatorname{GL}_{n}$ (via the natural projection $M\twoheadrightarrow\operatorname{GL}_{n}$).
@@
where the notation $\{a_{j}\}$ means: $K^{S}mK^{S}]$ acts as $\chi_{\operatorname{cycl}}(\operatorname{det}(m))^{-a_{j}}[K^{S}mK^{S}]$, where $m\in\operatorname{GL}_{n}(\mathbb{A}_{f}^{S})$, and for $u\in\mathbb{Z}_{p}^{\times}$, the operator $\langle u\rangle_{i}$ acts via $u^{a_{j}}\langle u\rangle_{i}$. This isomorphism is equivariant for the action of $U_{t}$ for $t\in T_{\operatorname{GL}_{n}}^{+}$ (recall that the action of these Hecke operators is automatically optimally normalised). This implies that the ordinary parts match up, and since (p1)(p-1) divides $a_{j}$, one also obtains an isomorphism after localising both sides at $\mathfrak{M}_{0}$. We therefore obtain an isomorphism
$\nu_{j}\colon\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})\xrightarrow{\sim}\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
@@
Let
$\varrho_{j}\colon G_{\mathbb{Q}}\to\operatorname{GL}_{n}\left(\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})/J_{j}\right)$
denote the semisimple Galois representation satisfying property (1) in the statement of the proposition, as constructed in [Sch15]. Set $\varrho^{\prime}_{j}=(\nu^{-1}_{j}\circ\varrho_{j})\otimes\omega^{-a_{j}}$, and $J_{j}^{\prime}=\nu^{-1}_{j}(J_{j})$; so the representation $\varrho^{\prime}_{j}$ is valued in $\operatorname{GL}_{n}(\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})/J_{j}^{\prime})$. Let $\chi_{\lambda(a_{j}),i}^{\prime}=\chi_{\lambda(a_{j}),i}\cdot\omega^{-a_{j}}$ valued in $\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})^{\times}$.
We now apply Proposition 9.5.8. More precisely, we obtain a $\mathcal{O}$-algebra morphism
@@
$\mathcal{T}_{w_{j}}\colon\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})\to\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
@@
$\mathcal{T}_{w_{j}}\circ\psi_{\widetilde{\lambda}_{j},n+1-i}^{-1}=\mathcal{T}_{w_{j}}\circ\psi_{\widetilde{\lambda}_{j},n+1+i}=\left\{\begin{array}[]{cc}\chi_{\lambda(a_{j}),w_{j}(i)}^{\prime}\otimes\omega^{-1}&\text{ if }1\leqslant w_{j}(i)\leqslant n\\ (\chi_{\lambda(a_{j}),2n+1-w_{j}(i)}^{\prime})^{-1}\otimes\omega&\text{ if }n+1\leqslant w_{j}(i)\leqslant 2n\end{array}\right$.
@@
As explained in Corollary 9.6.9 implies that there exists a continuous determinant
@@
which, after inverting $p$, is the determinant associated with a continuous semisimple Galois representation $\sigma\colon G_{\mathbb{Q}}\to\operatorname{GL}_{2n+1}(\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})[1/p])$ which satisfies:
  • For all $\ell\notin S$, the (reverse) characteristic polynomial of $\operatorname{Frob}_{\ell}^{-1}$ is the image of $\widetilde{H}_{\ell}(X)$.
  • One has
    $\sigma|_{G_{\mathbb{Q}_{p}}}\sim\left(\begin{array}[]{cccc}\psi_{\widetilde{\lambda}_{j},1}&*&*&*\\ &\psi_{\widetilde{\lambda}_{j},2}&*&*\\ &&\ddots&\vdots\\ &&&\psi_{\widetilde{\lambda}_{j},2n+1}\end{array}\right).$
Let $D=\mathcal{T}_{w_{j}}\circ\widetilde{D}$. Then $D$ modulo $J_{j}^{\prime}$ is the continuous determinant associated with $\varrho_{j}^{\prime}(-1)\oplus\mathbf{1}\oplus(\varrho^{\prime}_{j})^{\vee}(1)$. To ease notation, set
$\widetilde{A}_{0}$ $=\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})$
$A_{0}$ $=\mathbf{T}^{S,\operatorname{ord}}(\operatorname{H}^{j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
which are both semi-local finite $\mathcal{O}$-algebras. Let $A$ be a local direct factor of $A_{0}$ and let $\widetilde{A}$ be the corresponding local direct factor of $\widetilde{A}_{0}$. It suffices to work over these local algebras. Let $J$ denote the ideal $J^{\prime}_{j}$ in the local factor $A$. Since $D$ modulo $J$ is the continuous determinant associated with $\varrho_{j}^{\prime}(-1)\oplus\mathbf{1}\oplus(\varrho^{\prime}_{j})^{\vee}(1)$, we have
@@
where the overline denotes reduction modulo the maximal ideal of $A$.
@@
Our conditions on $\chi$ then imply that:
@@
We now argue as in the proof of [ACC+23, Proposition 4.4.6, Proposition 5.4.18]. More precisely, let $D_{A/J}=D\otimes_{A}A/J$, then we can identify
@@
corresponding to $\sigma$. From the local-global compatibility for $\sigma|_{G_{\mathbb{Q}_{p}}}$, we see that for elements $Y,Y_{1},\dots,Y_{2n+1}$ in $(A/J)[G_{\mathbb{Q}_{p}}]$, we have
@@
in $(A/J)[X]$ (note we may view $\psi_{\widetilde{\lambda}_{j},i}$ as valued in $A/J$), and
$(\prod_{i=1}^{2n+1}(\varrho^{\prime}_{j}(-1)(Y_{i})-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})),\prod_{i=1}^{2n+1}(\mathbf{1}(Y)-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})),\prod_{i=1}^{2n+1}((\varrho^{\prime}_{j})^{\vee}(1)(Y_{i})-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})))=(0,0,0)$ (9.6.17)
in $M_{n}(A/J)\times M_{1}(A/J)\times M_{n}(A/J)$. By choosing suitable idempotents as in the proof of [ACC+23, Proposition 5.4.18], we can deduce the desired properties for $\varrho^{\prime}_{j}$. More precisely, we can find an idempotent $e\in(A/J)[G_{\mathbb{Q}_{p}}]$ such that $\varrho^{\prime}_{j}(-1)(e)=1$ and the image of $e$ under the trivial representation and $(\varrho_{j}^{\prime})^{\vee}(1)$ is zero. We can also assume that $\psi_{\widetilde{\lambda}_{j},i}(e)=1$ if $\overline{\psi}_{\widetilde{\lambda}_{j},i}$ appears in $\overline{\varrho}^{\prime}_{j}(-1)$, and is zero otherwise.
Let $x,x_{1},\dots,x_{n}\in G_{\mathbb{Q}_{p}}$. Taking $Y=xe$ in the identity (9.6.16), we see that
@@
$(\prod_{k=1}^{n}(\varrho^{\prime}_{j}(-1)(x_{k})-\chi^{\prime}_{\lambda(a_{j}),k}\omega^{-1}(x_{k})),0,0)=(0,0,0).$

### v2

@@ 9.6.2. Some important lemmas
where $0\leqslant y<z\leqslant n+1$ are integers such that j=(y+1)(n+1)(y+1)y/2zj=(y+1)(n+1)-(y+1)y/2-z (if $y=0$, then ${\mathcal{B}}_{j}=\{z\}$, and if $z=n+1$, then ${\mathcal{B}}_{j}=\{1,\dots,y\}$). We set ${\mathcal{B}}_{0}=\varnothing$. Let $a\in\mathbb{Z}$ be an integer, and set $x_{j}=w_{{\mathcal{B}}_{j}}$ and $w_{j}=w_{{\mathcal{B}}_{j}^{c}}$.
@@
works (i.e., $\lambda=(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{i}=(2^{n}-2^{i})M$). Clearly part (1) of the lemma is satisfied for this weight, so we will check properties (2) and (3).
@@ 9.6.3. Mixed characteristic local-global compatibility
$\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})\cdot m\cdot\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})]\mapsto\overline{\psi}(\operatorname{det}m)\cdot\Theta([\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})\cdot m\cdot\operatorname{GL}_{n}(\widehat{\mathbb{Z}}^{S})]),\quad\quad m\in\operatorname{GL}_{n}(\mathbb{A}_{f}^{S}),$
@@
We first note that a Galois representation $\varrho_{j}$ satisfying (1) already exists from the work of Scholze [Sch15]. We therefore want to verify (2) and (3) for this representation. Since this only concerns the behaviour at $p$, we are free to enlarge $S$, increase $\mathcal{O}$ and deepen the level $K$ at primes in $S-\{p\}$ as we wish. Here we are using the fact that semisimple Galois representations over the Henselian semi-local $\mathcal{O}$-algebra $\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c,\lambda;j)$, reducing to $\overline{\rho}_{\mathfrak{m}_{0}}$ modulo each of its maximal ideals, are determined (up to isomorphism) by their associated determinant (see [Che14, Theorem 2.22]).
@@
Note $\operatorname{H}^{0}(X_{\operatorname{GL}_{1},C},\mathcal{O})$ is a finite free $\mathcal{O}$-module with rank equal to the number of connected components of $X_{\operatorname{GL}_{1},C}$, and contains the trivial representation $\mathcal{O}$ of the Hecke algebra for $\operatorname{GL}_{1}$. This implies that we have a natural map
@@
induced from mapping a Hecke operator for $M$ to one for $\operatorname{GL}_{n}$ (via the projection $M\twoheadrightarrow\operatorname{GL}_{n}$).
@@
of $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}$-modules, where the notation $\{a_{j}\}$ means the $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}$-action is twisted as follows:
  • for $m\in\operatorname{GL}_{n}(\mathbb{A}_{f}^{S})$, the symbol $K^{S}mK^{S}]$” in $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}$ acts as $\chi_{\operatorname{cycl}}(\operatorname{det}(m))^{-a_{j}}[K^{S}mK^{S}]$,
  • for $u\in\mathbb{Z}_{p}^{\times}$, the symbol $\langle u\rangle_{i}$” acts via $u^{a_{j}}\langle u\rangle_{i}$,
  • for $t\in T_{\operatorname{GL}_{n}}^{+}$, the operator $U_{t}$ acts as normal.
In particular, this isomorphism is $U_{t}$-equivariant (recalling that $U_{t}$ is automatically optimally normalised). This implies that the ordinary parts match up, and since (p1)(p-1) divides $a_{j}$, one also obtains an isomorphism after localising both sides at $\mathfrak{M}_{0}$. We therefore obtain an isomorphism
$\nu_{j}\colon\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})\xrightarrow{\sim}\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
@@
The ideal $\mathfrak{M}_{0}\subset\mathbf{T}_{\operatorname{GL}_{n}}^{S}$ satisfies the hypotheses of the proposition. Let
$\varrho_{j}\colon G_{\mathbb{Q}}\to\operatorname{GL}_{n}\left(\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})/J_{j}\right)$
denote the semisimple Galois representation attached to $\mathfrak{M}_{0}$ in [Sch15], which satisfies property (1) from the statement. Set $\varrho^{\prime}_{j}=(\nu^{-1}_{j}\circ\varrho_{j})\otimes\omega^{-a_{j}}$, and $J_{j}^{\prime}=\nu^{-1}_{j}(J_{j})$; so the representation $\varrho^{\prime}_{j}$ is valued in $\operatorname{GL}_{n}(\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})/J_{j}^{\prime})$. Let $\chi_{\lambda(a_{j}),i}^{\prime}=\chi_{\lambda(a_{j}),i}\cdot\omega^{-a_{j}}$ valued in $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})^{\times}$.
The twist by $\omega^{-a_{j}}$ ensures $\varrho_{j}^{\prime}$ satisfies the following local-global compatibility at $\ell\not\in S$: the (reverse) characteristic polynomial $\det(1-\operatorname{Frob}_{\ell}^{-1}X|\rho_{j}^{\prime})$ equals the image of the Hecke polynomial $H_{\ell}(X)$ in $(\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(\mathrm{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})/J_{j}^{\prime})[X]$.
Recall we defined
$\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})=\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\left(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}_{j},\mathcal{O}})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}\right).$
As $d-l(x_{j})=l(w_{j})=d-j$ (by Lemma 9.6.10), Proposition 9.5.8 (and its proof) yields an $\mathcal{O}$-algebra morphism
@@
$\mathcal{T}_{w_{j}}\colon\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})\to\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
@@
$\mathcal{T}_{w_{j}}\circ\psi_{\widetilde{\lambda}_{j},n+1-i}^{-1}=\mathcal{T}_{w_{j}}\circ\psi_{\widetilde{\lambda}_{j},n+1+i}=\left\{\begin{array}[]{cc}\chi_{\lambda(a_{j}),w_{j}(i)}^{\prime}\otimes\omega^{-1}&\text{ if }1\leqslant w_{j}(i)\leqslant n\\ (\chi_{\lambda(a_{j}),2n+1-w_{j}(i)}^{\prime})^{-1}\otimes\omega&\text{ if }n+1\leqslant w_{j}(i)\leqslant 2n\end{array}\right$. (9.6.14)
@@
Our set-up allows us to apply Corollary 9.6.9 to $\widetilde{\mathfrak{m}}$, giving a semisimple Galois representation
$\sigma_{\widetilde{\mathfrak{m}}}\colon G_{\mathbb{Q}}\to\operatorname{GL}_{2n+1}(\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})[1/p])$
satisfying local-global compatibility at $\ell\not\in S$ and at $p$, and a continuous determinant
@@
[…]
$(\mathcal{T}_{w_{j}}\circ\sigma_{\widetilde{\mathfrak{m}}}|_{G_{\mathbb{Q}_{p}}})^{\operatorname{ss}}=\mathbf{1}\oplus\bigoplus_{i=1}^{n}\chi^{\prime}_{\lambda(a_{j}),i}\omega^{-1}\oplus(\chi^{\prime}_{\lambda(a_{j}),i})^{-1}\omega,$ (9.6.15)
a representation $G_{\mathbb{Q}}\to\operatorname{GL}_{2n+1}(\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}}))$. Note this description of the semisimplification follows from (1b) of the corollary, and we thus deduce it takes integral values, even though $\mathcal{T}_{w_{j}}\circ\sigma_{\widetilde{\mathfrak{m}}}|_{G_{\mathbb{Q}_{p}}}$ itself is not necessarily integral.
By considering local-global compatibility at $\ell\not\in S$, we see $D$ modulo $J_{j}^{\prime}$ is the continuous determinant associated with $\varrho_{j}^{\prime}(-1)\oplus\mathbf{1}\oplus(\varrho^{\prime}_{j})^{\vee}(1)$. To ease notation, set
$\widetilde{A}_{0}$ $=\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(b,c,\widetilde{\lambda}_{j})=\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}_{j},\mathcal{O}})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}),$
$A_{0}$ $=\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}(\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda(a_{j}),\mathcal{O}})^{\operatorname{ord}}_{\mathfrak{M}_{0}})$
which are both semi-local finite $\mathcal{O}$-algebras. Let $A$ be a local direct factor of $A_{0}$ and let $\widetilde{A}$ be the corresponding local direct factor of $\widetilde{A}_{0}$. It suffices to work over these local algebras. Let $J$ denote the ideal $J^{\prime}_{j}$ in the local factor $A$; then $D$ modulo $J$ is the continuous determinant associated with $\varrho_{j}^{\prime}(-1)\oplus\mathbf{1}\oplus(\varrho^{\prime}_{j})^{\vee}(1)$. Let $\mathfrak{n}_{A}$ be the maximal ideal of $A$; then $D\hskip 2.0pt(\mathrm{mod}\hskip 2.0pt\mathfrak{n}_{A}):G_{\mathbb{Q}}\to\overline{\mathbb{F}}_{p}$ is the continuous determinant of $\overline{\varrho}_{j}^{\prime}(-1)\oplus\mathbf{1}\oplus(\overline{\varrho}^{\prime}_{j})^{\vee}(1)=\overline{\rho}_{\widetilde{\mathfrak{m}}}$. Its restriction to $G_{\mathbb{Q}_{p}}$ is also the continuous determinant associated to (9.6.15) modulo $\mathfrak{n}_{A}$. As semisimple Galois representations over $\overline{\mathbb{F}}_{p}$ are uniquely determined by their continuous determinant by [Che14, Thm. 2.12], we deduce
@@
where the overline denotes reduction modulo $\mathfrak{n}_{A}$.
@@
Our conditions on $\chi$ then imply that
@@
We now argue as in the proof of [ACC+23, Proposition 4.4.6, Proposition 5.4.18]. More precisely, let $D_{A/J}=D\otimes_{A}A/J$; then we can identify
@@
corresponding to $\sigma_{\widetilde{\mathfrak{m}}}$. From the local-global compatibility for $\sigma_{\widetilde{\mathfrak{m}}}|_{G_{\mathbb{Q}_{p}}}$, we see that for elements $Y,Y_{1},\dots,Y_{2n+1}$ in $(A/J)[G_{\mathbb{Q}_{p}}]$, we have
@@
in $(A/J)[X]$ (noting that, under $\mathcal{T}_{w_{j}}$, we may view $\psi_{\widetilde{\lambda}_{j},i}$ as valued in $A/J$), and
$\bigg(\prod_{i=1}^{2n+1}\Big(\varrho^{\prime}_{j}(-1)(Y_{i})-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})\Big),\prod_{i=1}^{2n+1}\Big(\mathbf{1}(Y)-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})\Big),\prod_{i=1}^{2n+1}\Big((\varrho^{\prime}_{j})^{\vee}(1)(Y_{i})-\psi_{\widetilde{\lambda}_{j},i}(Y_{i})\Big)\bigg)=(0,0,0)$ (9.6.19)
in $M_{n}(A/J)\times M_{1}(A/J)\times M_{n}(A/J)$. By choosing suitable idempotents as in the proof of [ACC+23, Proposition 5.4.18], we can deduce the desired properties for $\varrho^{\prime}_{j}$. More precisely, by (9.6.17) and our running assumptions, the representations $\varrho_{j}^{\prime}(-1)|_{G_{\mathbb{Q}_{p}}},\mathbf{1}$ and $(\varrho_{j}^{\prime})^{\vee}(1)|_{G_{\mathbb{Q}_{p}}}$ have no Jordan–Hölder factors in common, so we can find an idempotent $e\in(A/J)[G_{\mathbb{Q}_{p}}]$ such that $\varrho^{\prime}_{j}(-1)(e)=1$ and the image of $e$ under the trivial representation and $(\varrho_{j}^{\prime})^{\vee}(1)$ is zero. We can also assume that $\psi_{\widetilde{\lambda}_{j},i}(e)=1$ if $\overline{\psi}_{\widetilde{\lambda}_{j},i}$ appears in $\overline{\varrho}^{\prime}_{j}(-1)$, and is zero otherwise.
Let $x,x_{1},\dots,x_{n}\in G_{\mathbb{Q}_{p}}$. Taking $Y=xe$ in the identity (9.6.18), and using the description (9.6.14) of $\mathcal{T}_{w_{j}}\circ\psi_{\widetilde{\lambda}_{j},i}$, we see that
@@
$\bigg(\prod_{k=1}^{n}\Big(\varrho^{\prime}_{j}(-1)(x_{k})-\chi^{\prime}_{\lambda(a_{j}),k}\omega^{-1}(x_{k})\Big),0,0\bigg)=(0,0,0).$
Note that exactly these $n$ terms survive by (9.6.17).